Interest rates determine the present value of future amounts.​ (Round to the nearest​ dollar.

Requirements: 1. Determine the present value of five​-year bonds payable with face value of $ 91000 and stated interest rate of 10​%, paid semiannually. The market rate of interest is 10​% at issuance. 2. Same bonds payable as in Requirement​ 1, but the market interest rate is 16​%. 3. Same bonds payable as in Requirement​ 1, but the market interest rate is 8​%.
Answer #1

Solution 1:

Chart Values are based on:
n= (5 Years*2) 10 Half years
i= (10%/2) 5.00% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.613913 * $91,000 = $55,866
Interest (Annuity) [$91,000*10%*6/12] 7.721735 * $4,550 = $35,134
Price of Bonds $91,000

Solution 2:

Chart Values are based on:
n= (5 Years*2) 10 Half years
i= (16%/2) 8.00% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.463193 * $91,000 = $42,151
Interest (Annuity) [$91,000*10%*6/12] 6.710081 * $4,550 = $30,531
Price of Bonds $72,681

Solution 3:

Chart Values are based on:
n= (5 Years*2) 10 Half years
i= (8%/2) 4.00% Semi annual
Cash Flow Table Value * Amount = Present Value
Principal 0.675564 * $91,000 = $61,476
Interest (Annuity) [$91,000*10%*6/12] 8.110896 * $4,550 = $36,905
Price of Bonds $98,381